find B^1 for the five-point data set:

Question 4 1 pts Find , for the five-point data set: X: 226 8 10 y: 01233 01.29 O2.48 O 0.992 O 0.344 1.54 Question 5
Answer #1
X Y XY
2 0 0 4 0
2 1 2 4 1
6 2 12 36 4
8 3 24 64 9
10 3 30 100 9
Ʃx = 28
Ʃy = 9
Ʃxy = 68
Ʃx² = 208
Ʃy² = 23
Sample size, n = 5
x̅ = Ʃx/n = 28/5 = 5.6
y̅ = Ʃy/n = 9/5 = 1.8
SSxx = Ʃx² - (Ʃx)²/n = 208 - (28)²/5 = 51.2
SSyy = Ʃy² - (Ʃy)²/n = 23 - (9)²/5 = 6.8
SSxy = Ʃxy - (Ʃx)(Ʃy)/n = 68 - (28)(9)/5 = 17.6

Slope, b = SSxy/SSxx = 17.6/51.2 = 0.344

Answer: 0.344